OpenStudy (anonymous):

The price of a Ipod dropped from $299.99 to $180.55. What was the percent decrease in price

OpenStudy (anonymous):

math is my worst subject

mathslover (mathslover):

Ok : so given that : initial price : $299.99 final price : $180.55 . Right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol...

mathslover (mathslover):

So let us first find the decrease : can u find that ?

mathslover (mathslover):

decrease : initial price - final price

OpenStudy (anonymous):

39.18 % decrease.

OpenStudy (anonymous):

119.44

mathslover (mathslover):

Ok right carol :

OpenStudy (anonymous):

thx for the awnser but how doid u find it

OpenStudy (anonymous):

Sorry, I didn't realize mathslover was helping you already.

OpenStudy (anonymous):

and i have tor oudn the answer to the nearest hundreth

mathslover (mathslover):

Remember : percent decrease = =New amount-original amount/original amount X 100

OpenStudy (anonymous):

i have to round it to the nearest hundreth?

mathslover (mathslover):

So we have % decrease = decrease / initial price * 100

OpenStudy (lgbasallote):

it is traditional in mathematics word problems to add "Suppose" at the beginning of the word problem for completely improbable situations

mathslover (mathslover):

so % decrease = 119.44 / 299.99 * 100 % wht r u getting now ?

OpenStudy (anonymous):

oh ok now i get it. i am taking math online so im so lost on some of these questions

mathslover (mathslover):

Ok no problem we r here to help u \[\huge{\mathbb{Welcome}\textbf{To}\mathbb{Open}\textbf{Study}}\]

OpenStudy (anonymous):

Rounded to nearest hundredth percent makes it 39? or 39.80?

mathslover (mathslover):

39.80

OpenStudy (anonymous):

thast what i thought thx

mathslover (mathslover):

Gr8 .... congrats

OpenStudy (anonymous):

Joe Company in 2006 had sales of $800,000. In 2007 sales were up 75 percent. Calculate the sales for 2007

OpenStudy (anonymous):

this is the other one i had issues with the other 25 questions i figured out :)

mathslover (mathslover):

I have to go right now. . please post this as a new question .. . . some one will soon answer it